# Data-Driven Predictive Control for Robust Exoskeleton Locomotion

### v1

@@ II-A Models of Bipedal Locomotion
Full-Order Model: Bipedal locomotion can be described as a hybrid dynamic system consisting of domains with continuous dynamics and discrete transitions due to impact events [5]. In this work, we only consider a single domain with a single support phase where the stance foot is completely flat on the ground. We assume the ground height is always zero. The single support domain $\mathcal{D}$ and guard $\mathcal{S}$ of the system is defined as $\mathcal{D}=\{x\in T\mathcal{Q}:\textrm{p}_{\textrm{sw}}^{z}(q)\geq 0\}$ and $\mathcal{S}=\{x\in T\mathcal{Q}:\textrm{p}_{\textrm{sw}}^{z}(q)=0,\dot{\textrm{p}}^{z}_{\textrm{sw}}(q)<0\}$, where $\textrm{p}_{\textrm{sw}}(q)$ denotes the swing foot position, and the superscript denotes the axis we are considering (i.e., swing foot vertical position). The resulting hybrid system can be defined as follows:
$\mathcal{H}\mathcal{C}:$ $\dot{x}=f(x)+g(x)\,u$ $x\in\mathcal{D}\setminus\mathcal{S}$, (1)
$\mathcal{H}\mathcal{C}:$ $x^{+}=\Delta(x^{-})$ $x^{-}\in\mathcal{S}$, (2)
@@ II-B Behavioral Systems Theory
where $U_{p}\in\mathbb{R}^{\kappa T_{\textrm{ini}}\times(T-L+1)}$, $U_{f}\in\mathbb{R}^{\kappa N\times(T-L+1)}$, $Y_{f}\in\mathbb{R}^{\nu T_{\textrm{ini}}\times(T-L+1)}$,$Y_{f}\in\mathbb{R}^{\nu N\times(T-L+1)}$. From the Fundamental Lemma [18, 26, 27], any new trajectory lies in the range space of the Hankel matrices, or equivalently, there exists $\gamma\in\mathbb{R}^{T-L+1}$ such that
@@ III-A Construction of the Data-Driven Model
$\eta:=\textrm{col}(\textrm{p}_{\textrm{com2}\ell}^{x,y,z},\,\textrm{p}_{\textrm{com2r}}^{x,y,z})\quad\quad\mu:=\textrm{col}(\textrm{p}_{\textrm{cop2}\ell}^{x,y,z},\,\textrm{p}_{\textrm{cop2r}}^{x,y,z}).$
@@
Figure 5: Simulation comparison over nominal indicated by blue circles, DDPC indicated by orange stars, and MPC indicated by green diamonds. a) DDPC planner planning trajectories for increasing desired speed, capped at maximum step length 0.20.20.2 m at different step duration $t_{d}$. b) tracking performance over different desired step length with the same step duration. The dashed line indicate the ideal performance. Error bar indicates the standard deviation over 50 models. c) Simulation time before robot falling for the tracking performance comparison. Error bar indicates the standard deviation over 50 models. The maximum simulation time is 111111 s, indicated by the horizontal dash line. d) Comparison of Nominal and DDPC under time-varying perturbation applied on the negative $x$ direction.
@@ III-B DDPC Algorithm for Trajectory Optimization
$\mu_{k}\in\mathcal{U},\quad\eta_{k}\in\mathcal{H},\quad k=0,\ldots,N-1,$
where $\bm{r}^{\eta}:=\textrm{col}(r^{\eta}_{0},\cdots,r^{\eta}_{N-1})$ represents a reference output (i.e., CoM) trajectory, $\bm{r}^{\mu}:=\textrm{col}(r^{\mu}_{0},\cdots,r^{\mu}_{N-1})$ denotes a reference input (i.e., CoP) trajectory, and $(\bm{\mu}_{T_{\textrm{ini}}},\bm{\eta}_{T_{\textrm{ini}}})$ represents the pair of past actual I-O trajectories (i.e., feedback to the DDPC). In addition, $\mathcal{U}\subseteq\mathbb{R}^{4}$ and $\mathcal{H}\subseteq\mathbb{R}^{6}$ denotes the input and output feasibility sets, respectively. Finally, $Q\in\mathbb{R}^{6\times 6}$ and $R\in\mathbb{R}^{4\times 4}$ are chosen as positive definite weighting matrices.
@@ IV Layered Control Framework for Atalante
In this section, we introduce the remaining components and details for practically implementing the entire layered control framework used to realize locomotion on the lower-body exoskeleton Atalante (see Fig. 1). Atalante contains 12 actuated DoFs as illustrated in Fig. 3a, with three actuators on the hip, one on the knee, and two on the ankle for each leg. The device weighs 828282 (kg) and has adjustable thigh and shin length to account for the varying heights of users. Depending on the physical parameters of the users, the largest kinematically feasible step length in $x$ direction that the device is capable to achieve for flat-footed walking is less than 0.20.20.2 (m).
@@ IV-A Trajectory Planner with DDPC
We generate a reference trajectory with the open-source toolbox FROST [29] using the full-order system to account for kinematics and dynamics feasibility. This trajectory is described by 777-th order Bézier polynomials with the coefficient matrix of $\alpha=[\alpha_{\textrm{com}},\alpha_{\phi},\alpha_{\textrm{sw}}]$, where $\alpha_{\textrm{com}}$ describes the CoM trajectory, $\alpha_{\phi}$ described the pelvis orientation and $\alpha_{\textrm{sw}}$ described the swing foot position and orientation. The Bézier polynomials are evaluated based on a time-based phase variable. Specifically, considering a desired step duration $t_{d}$ and the initial time at the beginning of the domain $t_{0}$, we can calculate the phase variable $\tau_{k}=\frac{t_{k}-t_{0}}{t_{\textrm{d}}}$ at time point $t_{k}$. The reference for the CoM trajectory is then determined by $r^{\eta}_{k}=\textrm{col}(\textrm{p}_{\textrm{com}}^{x,y,z}(\alpha,\tau_{k})-p_{\ell}^{x,y,z}(\alpha,\tau_{k}),\,\,\textrm{p}_{\textrm{com}}^{x,y,z}(\alpha,\tau_{k})-\textrm{p}_{\textrm{r}}^{x,y,z}(\alpha,\tau_{k}))$. In addition, the reference CoP trajectory is generated via $r^{\mu}_{k}=\textrm{col}(-\textrm{p}_{\textrm{sw}}^{x,y}(\alpha,\tau_{k}),0,0)$ for the right stance and $r^{u}_{k}=\textrm{col}(0,0,-\textrm{p}_{\textrm{sw}}^{x,y}(\alpha,\tau_{k}))$ for the left stance.
Without explicit guidelines to construct the Hankel matrix for nonlinear systems, we empirically determine the hyperparameters of the DDPC algorithm via a grid search over the space of the discrete-time interval between nearby points $\delta_{t}\in[0.01,0.03]$, the trajectory length of $T\in[50,600]$, the initial trajectory length of $T_{\textrm{ini}}\in[5,50]$, and the control horizon of $N\in[10,300]$. This search space is constructed considering the computation speed, low-level controller frequency, noisy level of the data, and the amount of data required. We choose $T=400$, $T_{\textrm{ini}}=10$, $N=20$, and $\delta_{t}=0.02$ with a selection criteria on the accuracy of the least-square approximation over some unseen trajectory. In total, 8 (s) of data are used for the Hankel matrix construction, and a trajectory for 0.40.40.4 (s) is planned. This means our DDPC-based trajectory planner primarily acts as a short-term regulator to stabilize the system. We remark that even though $\textrm{p}_{\textrm{com}}^{z}$ is planned, it is not being used by the low-level controller but only used as part of the states to determine system dynamics. The trajectories are planned at 100 (Hz), which is faster than the interval $\delta_{t}$ specified. However, since the reference generation is based on the current domain and phase variable, the trajectory planner still has the ability to regulate the system behavior during this interval till the next time sample at which $\bm{\mu}_{T_{\textrm{ini}}}$ and $\bm{\eta}_{T_{\textrm{ini}}}$ are updated.
@@ IV-C Low-Level Feedback Controller
Our low-level controller, implemented in C++, receives trajectories for the CoM position target from the DDPC-based trajectory planner. Given that the controller operates at 1 (kHz), much faster than the planner’s replan frequency and the discrete-time interval, CoM target positions and velocities for each control tick is obtained via linear interpolation. Subsequently, we employ a Newton-Raphson numerical inverse kinematics (IK) algorithm to calculate the desired joint targets, which are then tracked using a passivity-based control method [30]. The kinematics and dynamics evaluation is performed with Pinocchio [31]. To account for uncertainty in state estimation and impact time, we switch to the following domain only when the swing foot ground reaction force exceeds some pre-defined threshold instead of time-based switching.
Figure 6: Gait tiles for simulation. Simulations with a) LIP-based MPC and b) DDPC controllers for walking at the speed of 0.160.160.16 (m/s), c) Nominal trajectory and d) DDPC under time-varying perturbations.
@@ V-A Simulation Results
We validate the effectiveness of our framework with numerical simulations in MuJoCo [33]. To account for uncertainty in mass distribution estimation for human users, we generated 50 randomized models with identical user total mass by varying the CoM offset and inertia properties, resulting in a $\textrm{p}_{\textrm{com}}$ ranging from [0.122,0.106][-0.122,-0.106][ - 0.122 , - 0.106 ] (m) in nominal standing configuration. The low level controller is using the information from the same nominal model constructed with parameters from [34].
@@
Tracking Performance: We compare tracking performance over these randomly generated models. For each model, simulation data with different desired step lengths ranging from 0.10.10.1 to 0.150.150.15 (m) were collected to construct the $\mathcal{G}$ matrix. An example trajectory generated by the DDPC planner and the corresponding actual CoM and CoP states are shown in Fig. 4. Moreover, DDPC is able to achieve stable walking in various speed as described in Fig. 5a. Additionally, we implemented an MPC planner based on the LIP dynamics for comparative analysis. This MPC shares a similar problem structure to that described in (III-B), but it substitutes the Hankel matrix trajectory constraints with those of the LIP dynamics. Considering the kinematic limits and feasibility concerns, we opted for a 0.80.80.8 (m) CoM z-position in the LIP model (as opposed to 0.850.850.85 m) and set $\delta_{t}=0.01$ (s) to discretize the LIP dynamics. The time horizon for the MPC is also chosen as $N_{\textrm{LIP}}=300$.
@@
The DDPC planner’s tracking performance was evaluated against that of the MPC and a nominal trajectory for desired walking speeds between 0.140.140.14 and 0.190.190.19 (m/s), with a constant step duration of 111 (s). As shown in Fig. 5b, the DDPC planner reliably achieves close tracking of the desired step length. As the desired speed increases by increasing the desired step length, the validity of LIP dynamics starts to degrade. In particular, the LIP MPC’s performance begins to decline at speeds above 0.160.160.16 (m/s), and the nominal trajectory cannot be appropriately tracked beyond 0.170.170.17 (m/s), as described in Fig. 5b. Furthermore, the DDPC planner maintains the system’s stability for a longer duration than the other controllers as demonstrated in Fig. 5c. Here, we define a robot maintaining CoM above a certain threshold throughout the maximum simulation time as a success. It is a failure if QP or IK becomes infeasible or CoM falls below a specified threshold. An example of failure case is shown in Fig. 6a for MPC. Since the DDPC-generated trajectory is based on the feasible trajectory used to construct $\mathcal{G}$, it tends to run into kinematics issues less frequently than physics-based reduced-order model like LIP, where no full-order model system information is exposed to the planner.
@@
Figure 8: The system evolution with the data-driven layered framework is shown in orange, and the system evolution with nominal trajectory is shown in blue. a) Experiment result for exoskeleton carrying 202020 (kg) of payload. CoP position in the foot frame for 5 left stance foot steps and 5 right stance foot steps. b) CoP position for exoskeleton with user inside. c) Gait tiles with the DDPC planner for the 202020 (kg) payload experiment. d) Gait tiles with the DDPC planner for the experiment with user.
@@
Adaptation to Perturbation: We further evaluate the DDPC’s robustness against time-varying external disturbances (see Fig. 6c and 6d). Specifically, we continuously apply an external perturbation force to the torso’s negative $x$ direction. This direction is chosen because the device is more sensitive towards perturbation applied along this axis with its fairly negative nominal CoM position. The initial perturbation force was set at 555 (N) with a discrete increase of 333 (N) every 333 (s). To adapt to this time-varying perturbation, the Hankel matrix is updated online. The choice of the Hankel matrix size requires a balance between satisfying persistently exciting requirements and using obsolete data. The simulation’s first 555 (s) were used for data collection, which was a smaller trajectory length $T=250$, and $\mathcal{G}$ is reconstructed every 1.51.51.5 (s). The least-square approximation step is completed within 7010070-10070 - 100 (ms). We ran the same setup for the aforementioned 50 randomly generated models and compared it with the nominal trajectory at 0.130.130.13 (m/s) desired speed. The DDPC has a higher success rate, defined by the percentage of the models being upright, compared to the nominal trajectory (see Fig. 5d).

### v2

@@ II-A Models of Bipedal Locomotion
Full-Order Model: Bipedal locomotion can be described as a hybrid dynamic system consisting of domains with continuous dynamics and discrete transitions due to impact events [5]. In this work, we only consider a single domain with a single support phase where the stance foot is completely flat on the ground. We assume the ground height is always zero. The single support domain $\mathcal{D}$ and guard $\mathcal{S}$ of the system is defined as $\mathcal{D}=\{x\in T\mathcal{Q}:\textrm{p}_{\textrm{sw}}^{z}(q)\geq 0\}$ and $\mathcal{S}=\{x\in T\mathcal{Q}:\textrm{p}_{\textrm{sw}}^{z}(q)=0,\dot{\textrm{p}}^{z}_{\textrm{sw}}(q)<0\}$, where $\textrm{p}_{\textrm{sw}}(q)$ denotes the swing foot position, and the superscript denotes the axis we are considering (i.e., swing foot vertical position). The resulting hybrid system $\mathcal{H}$ can be defined as follows:
$\mathcal{H}:$ $\dot{x}=f(x)+g(x)\,u$ $x\in\mathcal{D}\setminus\mathcal{S}$, (1)
$\mathcal{H}:$ $x^{+}=\Delta(x^{-})$ $x^{-}\in\mathcal{S}$, (2)
@@ II-B Behavioral Systems Theory
where $U_{p}\in\mathbb{R}^{\kappa T_{\textrm{ini}}\times(T-L+1)}$, $U_{f}\in\mathbb{R}^{\kappa N\times(T-L+1)}$, $Y_{p}\in\mathbb{R}^{\nu T_{\textrm{ini}}\times(T-L+1)}$,$Y_{f}\in\mathbb{R}^{\nu N\times(T-L+1)}$. From the Fundamental Lemma [18, 26, 27], any new trajectory lies in the range space of the Hankel matrices, or equivalently, there exists $\gamma\in\mathbb{R}^{T-L+1}$ such that
@@ III-A Construction of the Data-Driven Model
$\eta:=\textrm{col}(\textrm{p}_{\textrm{com2}\ell}^{x,y,z},\,\textrm{p}_{\textrm{com2r}}^{x,y,z})\quad\quad\mu:=\textrm{col}(\textrm{p}_{\textrm{cop2}\ell}^{x,y},\,\textrm{p}_{\textrm{cop2r}}^{x,y}).$
@@
Figure 5: Simulation comparison over nominal indicated by blue circles, DDPC indicated by orange stars, and MPC indicated by green diamonds. a) DDPC planner planning trajectories for increasing desired speed, capped at maximum step length $0.2$ m at different step duration $t_{d}$. b) tracking performance over different desired step length with the same step duration. The dashed line indicate the ideal performance. Error bar indicates the standard deviation over 50 models. c) Simulation time before robot falling for the tracking performance comparison. Error bar indicates the standard deviation over 50 models. The maximum simulation time is $11$ s, indicated by the horizontal dash line. d) Comparison of Nominal and DDPC under time-varying perturbation applied on the negative $x$ direction.
@@ III-B DDPC Algorithm for Trajectory Optimization
$\mu_{k}\in\mathcal{U},\quad\eta_{k}\in\mathcal{P},\quad k=0,\ldots,N-1,$
where $\bm{r}^{\eta}:=\textrm{col}(r^{\eta}_{0},\cdots,r^{\eta}_{N-1})$ represents a reference output (i.e., CoM) trajectory, $\bm{r}^{\mu}:=\textrm{col}(r^{\mu}_{0},\cdots,r^{\mu}_{N-1})$ denotes a reference input (i.e., CoP) trajectory, and $(\bm{\mu}_{T_{\textrm{ini}}},\bm{\eta}_{T_{\textrm{ini}}})$ represents the pair of past actual I-O trajectories (i.e., feedback to the DDPC). In addition, $\mathcal{U}\subseteq\mathbb{R}^{4}$ and $\mathcal{P}\subseteq\mathbb{R}^{6}$ denotes the input and output feasibility sets, respectively. Finally, $Q\in\mathbb{R}^{6\times 6}$ and $R\in\mathbb{R}^{4\times 4}$ are chosen as positive definite weighting matrices.
@@ IV Layered Control Framework for Atalante
In this section, we introduce the remaining components and details for practically implementing the entire layered control framework used to realize locomotion on the lower-body exoskeleton Atalante (see Fig. 1). Atalante contains 12 actuated DoFs as illustrated in Fig. 3a, with three actuators on the hip, one on the knee, and two on the ankle for each leg. The device weighs $82$ kg and has adjustable thigh and shin length to account for the varying heights of users. Depending on the physical parameters of the users, the largest kinematically feasible step length in $x$ direction that the device is capable to achieve for flat-footed walking is less than $0.2$ m.
@@ IV-A Trajectory Planner with DDPC
We generate a reference trajectory with the open-source toolbox FROST [29] using the full-order system to account for kinematics and dynamics feasibility. This trajectory is described by $7$-th order Bézier polynomials with the coefficient matrix of $\alpha=[\alpha_{\textrm{com}},\alpha_{\phi},\alpha_{\textrm{sw}}]$, where $\alpha_{\textrm{com}}$ describes the CoM trajectory, $\alpha_{\phi}$ described the pelvis orientation and $\alpha_{\textrm{sw}}$ described the swing foot position and orientation. The Bézier polynomials are evaluated based on a time-based phase variable. Specifically, considering a desired step duration $t_{d}$ and the initial time at the beginning of the domain $t_{0}$, we can calculate the phase variable $\tau_{k}=\frac{t_{k}-t_{0}}{t_{\textrm{d}}}$ at time point $t_{k}$. The reference for the CoM trajectory is then determined by $r^{\eta}_{k}=\textrm{col}(\textrm{p}_{\textrm{com}}^{x,y,z}(\alpha,\tau_{k})-p_{\ell}^{x,y,z}(\alpha,\tau_{k}),\,\,\textrm{p}_{\textrm{com}}^{x,y,z}(\alpha,\tau_{k})-\textrm{p}_{\textrm{r}}^{x,y,z}(\alpha,\tau_{k}))$. In addition, the reference CoP trajectory is generated via $r^{\mu}_{k}=\textrm{col}(-\textrm{p}_{\textrm{sw}}^{x,y}(\alpha,\tau_{k}),0,0)$ for the right stance and $r^{u}_{k}=\textrm{col}(0,0,-\textrm{p}_{\textrm{sw}}^{x,y}(\alpha,\tau_{k}))$ for the left stance.
Without explicit guidelines to construct the Hankel matrix for nonlinear systems, we empirically determine the hyperparameters of the DDPC algorithm via a grid search over the space of the discrete-time interval between nearby points $\delta_{t}\in[0.01,0.03]$, the trajectory length of $T\in[50,600]$, the initial trajectory length of $T_{\textrm{ini}}\in[5,50]$, and the control horizon of $N\in[10,300]$. This search space is constructed considering the computation speed, low-level controller frequency, noisy level of the data, and the amount of data required. We choose $T=400$, $T_{\textrm{ini}}=10$, $N=20$, and $\delta_{t}=0.02$ with a selection criteria on the accuracy of the least-square approximation over some unseen trajectory. In total, $8$ s of data are used for the Hankel matrix construction, and a trajectory for $0.4$ s is planned. This means our DDPC-based trajectory planner primarily acts as a short-term regulator to stabilize the system. We remark that even though $\textrm{p}_{\textrm{com}}^{z}$ is planned, it is not being used by the low-level controller but only used as part of the states to determine system dynamics. The trajectories are planned at $100$ Hz, which is faster than the interval $\delta_{t}$ specified. However, since the reference generation is based on the current domain and phase variable, the trajectory planner still has the ability to regulate the system behavior during this interval till the next time sample at which $\bm{\mu}_{T_{\textrm{ini}}}$ and $\bm{\eta}_{T_{\textrm{ini}}}$ are updated.
@@ IV-C Low-Level Feedback Controller
Our low-level controller, implemented in C++, receives trajectories for the CoM position target from the DDPC-based trajectory planner. Given that the controller operates at $1$ kHz, much faster than the planner’s replan frequency and the discrete-time interval, CoM target positions and velocities for each control tick is obtained via linear interpolation. Subsequently, we employ a Newton-Raphson numerical inverse kinematics (IK) algorithm to calculate the desired joint targets, which are then tracked using a passivity-based control method [30]. The kinematics and dynamics evaluation is performed with Pinocchio [31]. To account for uncertainty in state estimation and impact time, we switch to the next stance domain only when the swing foot ground reaction force exceeds some pre-defined threshold instead of time-based switching.
Figure 6: Gait tiles for simulation. Simulations with a) LIP-based MPC and b) DDPC controllers for walking at the speed of $0.16$ (m/s), c) Nominal trajectory and d) DDPC under time-varying perturbations.
@@ V-A Simulation Results
We validate the effectiveness of our framework with numerical simulations in MuJoCo [33]. To account for uncertainty in mass distribution estimation for human users, we generated 50 randomized models with identical user total mass by varying the CoM offset and inertia properties, resulting in a $\textrm{p}_{\textrm{com}}$ ranging from $[-0.122,-0.106]$ m in nominal standing configuration. The low level controller is using the information from the same nominal model constructed with parameters from [34].
@@
Tracking Performance: We compare tracking performance over these randomly generated models. For each model, simulation data with different desired step lengths ranging from $0.1$ to $0.15$ m were collected to construct the $\mathcal{G}$ matrix. An example trajectory generated by the DDPC planner and the corresponding actual CoM and CoP states are shown in Fig. 4. Moreover, DDPC is able to achieve stable walking in various speed as described in Fig. 5a. Additionally, we implemented an MPC planner based on the LIP dynamics for comparative analysis. This MPC shares a similar problem structure to that described in (III-B), but it substitutes the Hankel matrix trajectory constraints with those of the LIP dynamics. Considering the kinematic limits and feasibility concerns, we opted for a $0.8$ m CoM z-position in the LIP model (as opposed to $0.85$ m) and set $\delta_{t}=0.01$ s to discretize the LIP dynamics. The time horizon for the MPC is also chosen as $N_{\textrm{LIP}}=300$.
@@
The DDPC planner’s tracking performance was evaluated against that of the MPC and a nominal trajectory for desired walking speeds between $0.14$ and $0.19$ m/s, with a constant step duration of $1$ s. As shown in Fig. 5b, the DDPC planner reliably achieves close tracking of the desired step length. As the desired speed increases by increasing the desired step length, the validity of LIP dynamics starts to degrade. In particular, the LIP MPC’s performance begins to decline at speeds above $0.16$ m/s, and the nominal trajectory cannot be appropriately tracked beyond $0.17$ m/s, as described in Fig. 5b. Furthermore, the DDPC planner maintains the system’s stability for a longer duration than the other controllers as demonstrated in Fig. 5c. Here, we define a robot maintaining CoM above a certain threshold throughout the maximum simulation time as a success. It is a failure if QP or IK becomes infeasible or CoM falls below a specified threshold. An example of failure case is shown in Fig. 6a for MPC. Since the DDPC-generated trajectory is based on the feasible trajectory used to construct $\mathcal{G}$, it tends to run into kinematics issues less frequently than physics-based reduced-order model like LIP, where no full-order model system information is exposed to the planner.
@@
Figure 8: The system evolution with the data-driven layered framework is shown in orange, and the system evolution with nominal trajectory is shown in blue. a) Experiment result for exoskeleton carrying $20$ kg of payload. CoP position in the foot frame for 5 left stance foot steps and 5 right stance foot steps. b) CoP position for exoskeleton with user inside. c) Gait tiles with the DDPC planner for the $20$ kg payload experiment. d) Gait tiles with the DDPC planner for the experiment with user.
@@
Adaptation to Perturbation: We further evaluate the DDPC’s robustness against time-varying external disturbances (see Fig. 6c and 6d). Specifically, we continuously apply an external perturbation force to the torso’s negative $x$ direction. This direction is chosen because the device is more sensitive towards perturbation applied along this axis with its fairly negative nominal CoM position. The initial perturbation force was set at $5$ N with a discrete increase of $3$ N every $3$ s. To adapt to this time-varying perturbation, the Hankel matrix is updated online. The choice of the Hankel matrix size requires a balance between satisfying persistently exciting requirements and using obsolete data. The simulation’s first $5$ s were used for data collection, which was a smaller trajectory length $T=250$, and $\mathcal{G}$ is reconstructed every $1.5$ s. The least-square approximation step is completed within $70-100$ ms. We ran the same setup for the aforementioned 50 randomly generated models and compared it with the nominal trajectory at $0.13$ m/s desired speed. The DDPC has a higher success rate, defined by the percentage of the models being upright, compared to the nominal trajectory (see Fig. 5d).